THE WORKS OF PROFESSOR STANISŁAW GOŁAB IN THE FIELD OF FUNCTIONAL EQUATIONS

ROMAN GER

Abstract

Although the impressive scientific output of Professor Stanisław Gołab uniquely qualifies him as the very prominent Polish geometrist even an extremely brief (just one sentence altogether) biographic note in Nowa encyklopedia powszechna PWN, vol. 2, p. 567, states that S. Gołab’s papers deal also with functional equations. Actually, approximately 20% of his scientific publications might be classified among this theory. The list of 50 papers presented below (I will be frequently referring the listeners to this list during my talk) contains the items that are more or less directly related to functional equations. Among them the monograph [26] written jointly with János Aczél, the world-wide leader in this field, belongs to one of the first books devoted entirely to functional equations. The fundamental one among those considered in [26] is the translation equation

F (F (x, s), t) = F (x, ts)

intensively studied throughout several decades and nowadays. This equation, occurring while studying the problem of determination of geometric objects, yields also the subject of numerous papers listed below.

Another equation of basic importance is commonly known as the Gołab-Schinzel functional equation

f(x + yf(x)) = f(x)f(y)

which was obtained in [24] while looking for some special subgroups or subsemigroups of a given continuous group of transformations. Enormous number of mathematicians were and are inspired by that idea which led them to substantial generalizations of this equation studied in various classes of functions and pretty general spaces.

Equally inspiring was the paper [21] written jointly with Mieczysław Kucharzewski who, like his teacher S. Gołab, had great achievements in applying functional equation methods to geometry. Studying homomorphisms of the differential groups L_{n}^{s} and, in particular, the multiplicative Cauchy equation involving matrix arguments is extremely vivid up to now.

The other topics that will be discussed and reported on may be framed as follows:

• characterizations of mappings,

• other functional equations occurring in geometry,

• functional equations stemming from algebra,

• functional equations stemming from analysis,

• the significance of the domain of functional equation and the idea of local solutions,

• the functional equation of brigade;

plainly, all of them dealt with by Professor Stanisław Gołab in his papers presented below.

References

[1] Sur les fonctions homogènes, I. Équation d’Euler, C.R. Soc. Sci. Varsovie 25 (1932), 105-110.

[2] Sur la métrique angulaire des espaces généraux.I. Le pricipe de Chasles, Wiadom. Mat. 38 (1933), 1-12.

[3] Sur une définition axiomatique des nombres conjugués pour les nombres complexes ordinaires, Opusc. Math. (Cracov) 1 (1937), 1-11.

[4] Sur la possibilité de la détermination de “coniugum” dans le corps des nombres hypercomplexes à deux unités, Rev. Ci. (Lima) 40 (1937), 21-30.

[5] Sur quelques points concernant la notion du comitant, Ann. Soc. Polon. Math. 17 (1938), 177-192.

[6] Über die Klassifikation der geometrischen Objekte, Math. Z. 44 (1938), 104-114.

[7] Über eine Funktionalgleichung der Theorie der geometrischen Objekte, Wiadom. Mat. 45 (1938), 97-137.

[8] Sur la théorie des objets géométriques (Les objets diffèrentiels purs de deuxième et de troisième classe), Ann. Soc. Polon. Math. 19 (1946), 7-35.

[9] Sur la théorie des objets géométriques (Réduction de objets géométriques spéciaux de premiére classe aux objets du type Δ), Ann. Soc. Polon. Math. 20 (1947), 10-27.

[10] Sur la notion de dèrivée covariante, Colloq. Math. 1 (1948), 160.

[11] Sur les objets géométriques non-différentiels, Bull. Acad. Polon. Sci. Lettr. Cl. Sci. Math. Natur. Sér. A (1949), 67-72.

[12] Contribution à la théorie des objets géométriques, Prace Mat.-Fiz. (Warszawa) 47 (1949), 1-15.

[13] Sur les objets géométriques à une composante, Ann. Soc. Polon. Math. 23 (1950), 79-89.

[14] Sur la dèrivée covariante des objets géométriques de deuxième classe, Ann. Polon. Math. 1 (1954), 107-113.

[15] Über den Begriff der kovarianten Ableitung, Nieuw. Arch. Wisk. 2 (1954), 90-96.

[16] Géométrie différentielle vis-à-vis des hypothèses d’une faible régularité, Rev. Roumaine Math. Pures Appl. 1 (1956), 99-112.

[17] Zum distributiven Gesetz der reellen Zahlen, Studia Math. 15 (1956), 353-358.

[18] Un théorème sur la valeur moyenne θ dans la formule des accroissements finis, Ann. Polon. Math. 1 (1956), 118-125 (joint work with S. Łojasiewicz).

[19] Sur l’équation fonctionnelle f(X · Y) = f(X) · f(Y), Colloq. Math. 4 (1957), 265.

[20] Sur la l’algèbre des objets géométriques de première classe à une composante), Ann. Polon. Math. 4 (1957), 226-248 (joint work with H. Pidek).

[21] Über den Begriff der Pseudogröβen, Tensor N.S. 8 (1958), 78-89 (joint work with M. Kucharzewski).

[22] Sur l’équation f(X · Y) = f(X) · f(Y), Ann. Polon. Math. 6 (1959), 1-13.

[23] Zagadnienia teorii obiektów geometrycznych, Prace Mat. 3 (1959), 55-67.

[24] Sur l’équation fonctionnelle f[x + y · f(x)] = f(x) · f(y), Publ. Math. Debrecen 6 (1959), 113-125 (joint work with A. Schinzel).

[25] Sur les comitants algèbriques des tenseurs, Ann. Polon. Math. 9 (1960), 113-118.

[26] Funktionalgleichungen der Theorie der geometrischen Objekte, Monografie Mat. 39, PWN Warszawa, 1960 (a book written jointly with J. Aczél).

[27] On some new geometrical interpretations of the torsion of a skew curve, Ann. Polon. Math. 7 (1960), 269-278 (joint work with J. Kordylewski and M. Kuczma).

[28] Das Doppelverhältnis als Lösung einer Funktionalgleichung, Ann. Polon. Math. 9 (1960), 183-187 (joint work with J. Aczél, M. Kuczma and E. Siwek).

[29] Sur l’équivalence des objets géométriques de deuxième classe dont le nombre de composantes est ègal à la dimension de l’espace, Czechoslovak Math. J. 11 (1961), 475-479.

[30] La relation d’équivalence et les objets géométriques, Fund. Math. 50 (1962), 381-386.

[31] Über ein Funktionalgleichungssystem von O.E. Gheorgiu, Tagungsbericht “Funktionalgleichungen”, 7-11 Okt. 1963, Mathem. Forschungsinstitut Oberwolfach, pp. 9-10.

[32] Einige grundlegende Fragen der Theorie der Funktionalgleichungen, Glasnik Mat.- Fiz.. Astronom 20 (1965), 58-63.

[33] Über die Funktionalgleichung der Funktion Arccosinus, I. Die lokalen Lösungen, Publ. Math. Debrecen 12 (1965), 159-174 (joint work with L. Losonczi).

[34] La comparaison de la rapidité de convergence des approximations successives de la méthode de Newton avec la méthode de “regula falsi”, Mathematica (Cluj) 8 (1966), 45-49.

[35] Über ein System von Funktionalgleichungen, Ann. Polon. Math. 17 (1966), 223-243 (joint work with O.E. Gheorgiu).

[36] Über die Funktionalgleichung der Funktion Arccosinus, II, Publ. Math. Debrecen 13 (1966), 183-185 (joint work with L. Losonczi).

[37] Über Differentialkomitanten ersten Ordnung, I, Colloq. Math. 16 (1967), 173-184.

[38] Über Differentialkomitanten ersten Ordnung, II, Ann. Polon. Math. 20 (1968), 81-89.

[39] Über die Grundlagen der affinen Geometrie, Jber. Deutsch. Math.-Verein. 71 (1969), 138-155.

[40] Über ein Funktionalgleichungssystem aus dem Gebiete der Geometrie, I Intern. Kolloquium über Funktionalgleichungen und ihre Anwendungen, Publ. of the Technical University, Miskolc (1970), 237-242.

[41] Functional equations in geometry, Zeszyty Nauk. Uniw. Jagiellońskiego. Prace Mat. 14 (1970), 13-19.

[42] Remarks on one-parameter subsemigroupes of the affine group and their homo- and isomorphisms, Aequationes Math. 4 (1970), 1-10 (joint work with J. Aczél).

[43] Über das Flächenmass in Vektorräumen, Prace Mat. 14 (1970), 51-57 (joint work with M. Kucharzewski).

[44] Les équations fonctionnelles et la théorie des groupes, Rev. Roumaine Math. Pures Appl. 15 (1970), 1395-1406 (joint work with Z. Krzeszowiak).

[45] Sur l’équation fonctionnelle des brigades, [Corso tenuto a La Mendola (Trento) dal 20 al 28 agosto 1970], Edizioni Cremonese, Roma, 1971, pp. 143-151.

[46] Note on inner product in vector spaces, Aequationes Math. 8 (1972), 74-75 (joint work with H. Światak).

[47] Ein Beitrag zur Axiomatisierung des Begriffs des Skalarproduktes, Demonstratio Math. 5 (1973), 273-288.

[48] Sur un systéme d’équations fonctionnelles lié au rapport anharmonique, Ann. Polon. Math. 29 (1974), 273-280.

[49] Theorie der Quasigruppen, [In: Grundlagen der Geometrie und algebraische Methoden, Internat. Kolloq., Pädagog. Hochsch. “Karl Liebknecht”, Potsdam, 1973], Potsdamer Forschungen, Reihe B, Heft 3, Potsdam, 1974, pp. 137-151.

[50] Über das Carnotsche Skalarprodukt in schwach normierten Vektorräumen, Aequationes Math. 13 (1975), 9-13.

Institute of Mathematics Silesian University Bankowa 14, 40-007 Katowice POLAND

http://www.imi.wsp.czest.pl/pracownicy/ger.html

e-mail: romanger@us.edu.pl